OpenStudy (anonymous):

what is the simplified form of ((x)/(y^2))/((x^2)/(y^5))? Can anyone solve this problem?

OpenStudy (anonymous):

multiply by the reciprocal i.e. write \[\frac{x}{y^2}\times \frac{y^5}{x^2}\] then cancel

OpenStudy (anonymous):

okay could you please draw it so i could see it please

OpenStudy (anonymous):

@@satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

what is \(\frac{x}{x^2}\) ?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

i mean (1)/(x) @satellite73

OpenStudy (anonymous):

@ghazi mathlegend said you could probably solve this

OpenStudy (anonymous):

yes it is \(\frac{1}{x}\)

OpenStudy (anonymous):

how about \(\frac{y^5}{y^2}\) ?

OpenStudy (anonymous):

What is this about?

OpenStudy (anonymous):

because \[\frac{xy^5}{x^2y^2}=\frac{x}{x^2}\times \frac{y^5}{y^2}\]\]

OpenStudy (anonymous):

in other words, you get \(\frac{1}{x}\times y^3=\frac{y^3}{x}\)

OpenStudy (anonymous):

ohh that makes sense lol thankyou.